OpenStudy (anonymous):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. If Michelle returned the movie and paid $4.00, how many nights did she keep the movie? 1 2 3 4

OpenStudy (anonymous):

need help n explain

OpenStudy (anonymous):

1.25x + 1.5 = 4 1.25x = 4 - 1.5 1.25x =2.5 x = 2 nights.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

ur welcome